The exercise price of the options is $100 per share, all options are European and the...

  1. The exercise price of the options is $100 per share, all options are European and the stock does not pay any dividend. The call price is $25 per share and the put price is $17 per share. Both options mature in 2 years. Finally, the annual rate of interest is 5% with continuous compounding.   What is the stock price today?
Answer #1

Using Put call parity
S+P=C+Xe^(-rt)
=>S=C+Xe^(-rt)-P
=>S=25+100*e^(-5%*2)-17
=>S=98.4837418